9. Application - Elasticity and hotel rooms 


The following graph input tool shows the daily demand for hotel rooms at the Triple Sevens Hotel and Casino in Las Vegas, Nevada. To help the hotel management better understand the market, an economist identified three primary factors that affect the demand for rooms each night. These demand factors, along with the values corresponding to the initial demand curve, are shown in the following table and alongside the graph input tool. 
Answer #1

If average household income increases by 50%, from $40,000 to $60,000 per year, the quantity of rooms demanded at the Triple Sevens ( rises)as people have more to spend from rooms 150 per night to 200 rooms per night. Therefore, the income elasticity of demand is (positive), as when income rises demand rises, meaning that hotel rooms at the Triple Sevens are (a normal good)

If the price of a room at the Exhilaration were to decrease by 10%, from $250 to $225, while all other demand factors remain at their initial values, the quantity of rooms demanded at the Triple Sevens (falls, as people would shift there) from 150 rooms per night to 145 rooms per night. Because the cross-price elasticity of demand is (positive, as when Exhilaration price fell demand for Triple Sevens also fell), hotel rooms at the Triple Sevens and hotel rooms at the Exhilaration are (substitutes)

Triple Sevens is debating decreasing the price of its rooms to $325 per night. Under the initial demand conditions, you can see that this would cause its total revenue to (increase, earlier revenue = 150*350 = 52500, now revenue = 175*325 = 56875). Decreasing the price will always have this effect on revenue when triple sevens is operating on the (elastic, where same quantity is preferred at different prices) portion of its demand curve.